OpenStudy (anonymous):

x/6 = x/8+5 How do you start off solving this equation?

OpenStudy (anonymous):

maybe multiply both sides by 24 to clear the annoying factions

OpenStudy (blurbendy):

I would start by getting the x fractions together on the same side. Then add them.

OpenStudy (anonymous):

So i would divide the fractions to get them all on one side?

OpenStudy (blurbendy):

I would recommend moving the x fraction on the right, to the left side. You can do this simply by subtracting the whole expression. No need to divide.

OpenStudy (blurbendy):

Do you follow?

OpenStudy (anonymous):

Okay i kind of understand

OpenStudy (blurbendy):

So what does the equation look like after you subtract x / 8 to the right side?

OpenStudy (blurbendy):

i mean to the left side, sorry

OpenStudy (anonymous):

i got x/2+5

OpenStudy (blurbendy):

The equation should look like this: x / 6 - x / 8 = 5

OpenStudy (anonymous):

I was right the first time ugh!!

OpenStudy (blurbendy):

Now, we need a common denominator so that we can subtract them. The least common denominator is 24, so we can rewrite the equation as: 4x / 24 - 3x /24 = 5 and now we can subtract them. So what does the equation look like after the subtraction?

OpenStudy (anonymous):

Ok hold on

OpenStudy (anonymous):

x/24=5

OpenStudy (blurbendy):

yup! almost done. we need x by itself, so we can multiply both sides by 24. What does the equation look like after we multiply both sides by 24?

OpenStudy (anonymous):

x=120?

OpenStudy (blurbendy):

Perfect. Good job!

OpenStudy (anonymous):

Wow thanks